OpenStudy (jamaljunior):

Solve for d. 6(d+1)−2d=54

OpenStudy (vuriffy):

6(d+1)−2d=54 6d + 6 - 2d = 54 Combine like-terms 4d + 6 = 54 -6 4d = 48 48/4 = d What does d = ?

OpenStudy (jamaljunior):

12

OpenStudy (vuriffy):

Correct!